OpenStudy (anonymous):

A field is to be fertilized at a cost of $0.08 per square yd. The rectangular part of the field is 95 Yd Long and the two Semi Circles are 49 Yd long. Find the cost of fertilizing the field. DO NOT GIVE ME THE ANSWER HELP ME SO I CAN DO FUTURE PROBLEMS.

OpenStudy (anonymous):

1st find the total area in term of sq.yd ...you know how ..???

OpenStudy (anonymous):

Uh Hold on.

OpenStudy (anonymous):

@zaibali.qasmi Do you add both of the problems together or do you times them?

OpenStudy (anonymous):

1st find separatly and then add them up to get the total..!!!

OpenStudy (anonymous):

Sorry I mean 193

OpenStudy (anonymous):



OpenStudy (anonymous):

Uhm I got 358109.15

OpenStudy (anonymous):

The field area can also be found by adding the area of the circle to the rectangle's area.

OpenStudy (anonymous):

try again..!!!its not accurate..

OpenStudy (anonymous):

Ok But do I use the area of the rectangle as R?

OpenStudy (anonymous):

I am so confused now.

OpenStudy (anonymous):

\[\pi \left(\frac{49}{2}\right)^2+49\text{*} 95=6541 \text{ rounded} \]

OpenStudy (anonymous):

Oh THANKYOU THANKYOU~!!!!

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

But Now what do I need to do?

OpenStudy (anonymous):

yes..now you have given that 0.08 dollar per sq.yd is expenditure..now just multiply the area with 0.08 to find total xpenditures for whole field..you will get result.

OpenStudy (anonymous):

Ok Now I get it So the answer is $523.18

OpenStudy (anonymous):

So for the formula\[\pi * SemiCircLength \div NumberOfCemiCircles + SemiCircLength * SizeOfSquare = Answer\]

OpenStudy (anonymous):

no...\[{\pi*(semi circle lenght)^2}/2+{\pi*(semi circle lenght)^2}/2+ area of rectangle\]